OpenStudy (anonymous):

The activation energy of a certain reaction is 33.3 kJ/mol . At 24 degrees C , the rate constant is 0.0110s^-1 . At what temperature in degrees Celsius would this reaction go twice as fast?

OpenStudy (anonymous):

Use the Arrhenius equation \[\ln \frac{ k2 }{ k1 }=-\frac{ Ea }{ R }(\frac{ 1 }{ T2 }-\frac{ 1 }{ T1 }\]